# Magnetoconvection in a spherical shell: Equatorial symmetry during the transition from the weak- to the strong-field regime

### 09 Mar 2026 v1

@@ 3.2 Dynamo models
Figure 2: The $(l,m)=(1,0)$ component of the poloidal scalar, $p_{10}$, at the outer boundary, for dynamo simulations with varied $\rm Pm$} and $\rm Ra$$^{\prime}$ close to the onset of (hydrodynamic) convection. (This data is derived from simulations conducted by Teed and Dormy (2025b, a)).
@@
which is overlaid as a dashed line on Figure 2. This is too weak to significantly alter the morphology of the flow, although it can adjust the preferred mode, as in Figure 3(a), 3(b), which are dominated by $m=6$ and $m=4$ modes, respectively, as opposed to the preferred hydrodynamic onset mode of $m=7$. This saturation amplitude is almost independent of $\rm Pm$}, whilst dynamo onset is not, and so, whilst dynamo onset can occur almost arbitrarily close to ${\rm Ra}^{\prime}=1$ by increasing $\rm Pm$}, the saturated magnetic field strength should tend to zero in that limit. However, it has also been hypothesised that a subcritical dynamo, in which dynamo action occurs at Rayleigh numbers below hydrodynamic onset, could be possible since the onset of convection can be decreased by a sufficiently large magnetic field. Fearn (1979) predicts that for asymptotically small $\rm Ek$}, $p_{10}\sim B_{0}\gtrsim\mathcal{O}\left({\rm Ek}^{1/6}\right)$ should be large enough to begin to decrease ${\rm Ra}_{c}$. Mason et al. (2022) find that ${\rm Ek}=10^{-4}$ is not sufficiently small to be asymptotic and that a larger field, $B_{0}=\mathcal{O}\left(1\right)$ is required to alter ${\rm Ra}_{c}$ (and also depends on $\rm Pm$}). These scalings are not amenable to the search for a subcritical dynamo.
@@
As the Rayleigh number increases (${\rm Ra}^{\prime}\gtrsim 5$), inertia grows and convective rolls can become unstable to centrifugal instabilities. This characterises the onset of inertia-dominated solutions; these solutions can also sustain a magnetic field (for sufficiently large $\rm Pm$}), although they have a tendency to be multipolar rather than dipolar due to the smaller inertial(/Coriolis) scales (compared with the (VAC) onset scales). However, the length scales on the strong-field dipolar (MAC) branch are larger since the magnetic field can overcome the resistance of large length scales perpendicular to the rotation axis from the Coriolis effect. The larger length scales on this branch reduce the energy lost to diffusion and hence increase the strength of the dipolar field; this feedback mechanism is another potential avenue for bistability.
The energy in zonal flow, which is on par with the other two components in hydrodynamic convection, is found to drop significantly in dynamo simulations, since it is resisted by the presence of a large scale poloidal field (Aubert, 2005). We investigate this aspect of the problem in §4, in our range of magnetoconvection simulations. In the absence of a significant $\Omega$-effect from a zonal flow, the dynamo mechanism on all three branches must be of an $\alpha^{2}$-type. A straightforward balance shows that the toroidal component of the field should therefore be of comparable strength to the poloidal field (Roberts and Soward, 1992). Teed and Dormy (2025b) verify this in the strong-field regime, but find that the toroidal field can be several times larger in the weak-field regime; a feature that can perhaps be explained by the disparity between the strengths of the axial and columnar flows in that case.
@@
Figure 4 shows the columnarity and symmetry in dynamo simulations for $\rm Ra$$^{\prime}$ between the onset of dynamo action and ${\rm Ra}^{\prime}=5$ for the two values of $\rm Pm$} that exhibit the clearest bistability between weak- and strong-field (dipolar) solutions. The sharp drops in equatorial symmetry at ${\rm Ra}^{\prime}=3.3$, for ${\rm Pm}=5$, and ${\rm Ra}^{\prime}=2.1$, for ${\rm Pm}=12$, indicate that an important element in the onset of strong-field convection is the breaking of equatorial symmetry which, as in hydrodynamic convection, seems to occur due to the destabilisation of antisymmetric modes of convection. Polar convective modes are significantly destabilised by the presence of an axial magnetic field since the wavelength of the onset mode perpendicular to the rotation axis is short, minimising the effect of rotation, which is felt more strongly inside the tangent cylinder; this also corresponds to the form required to minimise the bending of axial field lines (Sakuraba, 2002). The columnarity drops at the onset of the polar modes due to the growth of the magnetic field which can then destabilise convective rolls in the flow outside the tangent cylinder.
@@ 4.1 Magnetoconvection with ${\rm Pm}=1$
Figure 6: Time-averages of the compensated-curl of forces in simulations at ${\rm Pm}=1$, $p_{10}=0.1,1.0$, at three values of $\rm Ra$$^{\prime}$. Odd and even $l$-modes are presented separately in the first two cases due to the strong equatorial symmetry of the buoyancy force and the prominence of the $m=8$ convective mode and $m=0$ zonal mode leading to a dearth of odd-$l$ buoyancy modes.
@@ 4.2 Magnetoconvection with ${\rm Pm}=5$
Figure 10: Time-averaged force spectra as in Figure 6, at ${\rm Pm}=5$, for ${\rm Ra}^{\prime}=2.5$ and $5.0$ for $p_{10}=0.1$ and ${\rm Ra}^{\prime}=1.6,1.8,2.5$, and $5.0$ for $p_{10}=1.0$.
@@
Figures 10(a),10(b), and10(d) show the force balances at ${\rm Ra}^{\prime}=1.6,1.8,2.5$, below, within, and above the vacillating regime, respectively. In each case the force balance could be described as MAC, as the Lorentz force is consistently part of the dominant balance across a range of scales, however it is clear that there are subtle distinctions between the three cases. At ${\rm Ra}^{\prime}=1.6$ and ${\rm Ra}^{\prime}=1.8$, the Lorentz force does not enter the dominant balance at the largest scales, entering sharply at $l=8$ for ${\rm Ra}^{\prime}=1.6$, when the columns are steady, and gradually at $l\approx 10$, for ${\rm Ra}^{\prime}=1.8$, when the columns are vacillating. The remaining forces are also smoother at ${\rm Ra}^{\prime}=1.8$ than at ${\rm Ra}^{\prime}=1.6$, consistent with the wide range of scales that emerge due to the break-up of the columns. Above the vacillating regime the Lorentz force also enters the dominant balance at smallest $l$, and viscosity and inertia are increasing subdominant.
@@ 4.3 Magnetoconvection with ${\rm Pm}=12$
Figure 19(a) shows the zonally averaged poloidal streamlines for a weak-field solution at a bistable point in parameter space. The streamlines are equatorially symmetric and primarily reflect the axial circulation within convection columns, with neighbouring counter-rotating cells persisting after zonal averaging due to slight asymmetries between cyclonic and anti-cyclonic columns. On the strong-field branch, the zonally averaged streamlines remain largely equatorially symmetric (Figure 19(d)), consistent with $s_{u}>0$ (Figure 5(f)), but the circulation cells extend much further into the equatorial region and inside the tangent cylinder. This reflects the increased magnetic field strength, which allows the Lorentz force to overcome the rotational constraint in regions where the boundaries are closest and (hence) the Coriolis force is most strongly felt.
@@ 5 Conclusions
In §3.2 we find that the weak-strong transition in dynamo simulations leads to a spontaneous breaking of the equatorial symmetry of the weak-field, highly-columnar, state, as the Rayleigh number is increased. There are three plausible candidates for the onset mechanism of the antisymmetric modes:
@@
    That a new instability occurs, possibly triggered by, but evolving independently of, the existing convection rolls.
Case (i) seems plausible since the inertia of convection rolls increases with the Rayleigh number. At sufficiently large $\rm Ra$} columnar flow can then become unstable to a buckling instability (like the Billant instability), which could lead to the emergence of modes antisymmetric about the equator. This hypothesis, however, is not supported by §3.1, and Figures 1(a),1(b), in particular, which show that the columnarity of the flow (in the bulk of the shell) is reduced at lower Rayleigh numbers than the occurrence of symmetry-breaking, which occurs at ${\rm Ra}^{\prime}\approx 10$. By this point the columnarity of the flow has fallen to around $0.6$, which is significantly smaller than in some (weakly supercritical) strong-field states (Figures 4(a),4(b)). Though this instability may be responsible for an alternative transition (at these large Ra) to inertially-driven/inertially-influenced states that may correspond/be relevant to multipolar solutions commonly found in dynamo simulations.
@@
Gilman (1975) and Sreenivasan and Jones (2006) have previously considered the consequences of the onset of convection inside the tangent cylinder, which can be well-approximated by the linear instability of a magnetic conduction state due to the relative quiescence of the tangent cylinder as a result the enhanced constraint of rapid rotation due to the closeness, in $z$, of the inner and outer boundaries. The Lorentz force, in rapidly-rotating magnetoconvection, can overcome the Taylor-Proudman constraint at significantly lower Rayleigh numbers than the viscous force alone (Fearn, 1979). Sreenivasan and Jones (2006) find that an axial magnetic field reduces the critical Rayleigh number by around a factor of three, and the further addition of a strong toroidal magnetic field, which is also present in the tangent cylinder in models of weak-field convection (Figure 14), would only compound this effect. At ${\rm Pm}=5$, symmetry-breaking can occur at around ${\rm Ra}^{\prime}=3.3$, (three times less than in purely hydrodynamic simulations where the symmetry-breaking must wait for inertial effects), whilst at ${\rm Pm}=12$, this occurs at ${\rm Ra}^{\prime}=2$.
Appendix B outlines a possible reasoning for polar convection ti appear preferentially as an antisymmetric, rather than an symmetric mode , despite the disconnection of the regions within the tangent cylinder above and below the inner core. Hence, it seems that by measuring the departure of the flow from equatorial symmetry, we do indeed capture the onset of polar convection.
Untangling the cause and effect in the correlation between the onset of polar convection and the onset of the strong-field regime is challenging and could be an interesting subject for future work. Skene et al. (2024) investigate the case of bistability between a hydrodynamic and a weak-field state using an adjoint method, and hence are able to determine an unstable fixed point that influences the evolution of an initial state towards either a successful or failed dynamo. If similar methods can be employed between the weak- and strong-field branches it may help to understand the exact role that the onset of polar convection has in the development of the strong-field regime.
@@ Appendix A Magnetic boundary conditions
A.1 Electrically-conducting boundaries
To a reasonable approximation, the Earth’s mantle is electrically insulating. In an insulating medium,
@@ A.1 Electrically-conducting boundaries
Decomposed into poloidal and toroidal components, $\boldsymbol{B}$ and $\boldsymbol{j}$ are
@@
$\boldsymbol{j}$ $=\boldsymbol{\nabla}\times\boldsymbol{\nabla}\times\left(j_{p}\boldsymbol{r}\right)+\boldsymbol{\nabla}\times\left(j_{t}\boldsymbol{r}\right).$ (26b)
Which, by definition, are related by
@@
hence, the magnetic field in an insulating medium must satisfy the
@@
$B_{p}=\sum_{l}\sum_{m}P^{m}_{l}\left(r\right)Y^{m}_{l}\left(\theta,\phi\right),\quad B_{t}=\sum_{l}\sum_{m}T^{m}_{l}\left(r\right)Y^{m}_{l}\left(\theta,\phi\right)$ (29)
@@
If there are no external sources then the exterior magnetic field is $P^{m}_{l}=Ar^{-l-1}$, for some constant $A$, and so the boundary condition on the internal field, matching with this solution is
$\left(r\frac{{\rm d}}{{\rm d}r}+l\left(l+1\right)\right)P^{m}_{l}\biggr|_{r=r_{O}}=0,\quad\forall\penalty 10000\ l,m.$ (31)
At the inner boundary the role of internal and external sources are reversed by the geometry in relation to increasing $r$, i.e., sources correspond to $\alpha=l$, hence, the condition at the inner boundary is
@@
Say an external source exists which induces a uniform axial magnetic field of the form $\boldsymbol{B}=B_{0}\boldsymbol{\hat{z}}$ where $\boldsymbol{\hat{z}}=\cos{\theta}\boldsymbol{\hat{r}}-\sin{\theta}\boldsymbol{\hat{\theta}}$. Then, given $B_{r}=r^{-1}L^{2}B_{p}$ (where $L^{2}=l(l+1)$), this axial field takes the form
$B_{p}(r,\theta,\phi)=P^{0}_{1}(r)\sqrt{3}\cos{\theta},\quad P^{0}_{1}(r)=\frac{B_{0}}{2\sqrt{3}}r.$ (33)
@@
$P^{0}_{1}=\frac{B_{0}}{2\sqrt{3}}r+\frac{A^{0}_{1}}{r^{2}},\qquad P^{m}_{l}=A^{m}_{l}r^{-l-1},\penalty 10000\ \left(l,m\right)\neq\left(1,0\right),$ (34)
@@
$\left(\frac{{\rm d}}{{\rm d}r}+\frac{2}{r}\right)B_{P}\biggr|_{r=r_{O},\penalty 10000\ (l,m)=(1,0)}=\frac{\sqrt{3}B_{0}}{2},$ (35)
@@
A.2 Perfectly-conducting boundaries
In magnetoconvection problems, an external magnetic field is applied to the system, that is, we can write
$\boldsymbol{B}=\boldsymbol{B}_{0}+\boldsymbol{B}^{\prime},$ (36)
where $\boldsymbol{B}_{0}$ is a fixed, known field, and $\boldsymbol{B}^{\prime}$ is induced by the motion of the fluid. Here, as in many other studies, we assume that the imposed field is axial and uniform, i.e.,
$\boldsymbol{B}_{0}=B_{0}\boldsymbol{\hat{z}}$. (37)
One option now is to assume that the varying field, $\boldsymbol{B}^{\prime}$, is insulating
Another option is to fix the magnetic field at the outer boundary. By this we mean that $B_{P}\left(r_{\rm o},l,m\right)$ is fixed for all $l,m$ (generally, $B_{P}(r_{\rm o},l,m)=0$ for all $(l,m)\neq(1,0)$). Immediately, this is different to (31) and (34), since this condition applies directly to $B_{P}$ rather than its derivative.
In the GSD code, say we set $P_{1}^{0}(r_{\rm o})=C$. The radial component of the magnetic field is given by
@@ A.2 Perfectly-conducting boundaries
and hence, setting all $P^{m}_{l}=0$, $\left(l,m\right)\neq\left(1,0\right)$,
$B_{r}\left(r_{\rm o},\theta,\phi\right)=2\frac{C}{r_{\rm o}}\left[\frac{1}{2}\frac{{\rm d}}{{\rm d}\mu}\left(\mu^{2}-1\right)\right]_{\mu=\cos{\theta}}=\frac{2C}{r_{\rm o}}\cos{\theta}=\frac{13C}{10}\cos{\theta},$ (39)
where $r_{\rm o}=20/13$. This can be easily verified. Furthermore, the appropriate scale factor between the external field, $B_{0}$, and the fixed poloidal, $C$, boundary conditions is $2/r_{\rm o}$.

### 17 Mar 2026 v2

@@ 3.2 Dynamo models
Figure 2: The $(l,m)=(1,0)$ component of the poloidal scalar, $p_{10}$, at the outer boundary, for dynamo simulations with varied $\rm Pm$} and $\rm Ra$$^{\prime}$ close to the onset of (hydrodynamic) convection. (This data is derived from simulations conducted by Teed and Dormy, 2025a, b).
@@
which is overlaid as a dashed line on Figure 2. This is too weak to significantly alter the morphology of the flow, although it can adjust the preferred mode, as in Figure 3(a), 3(b), which are dominated by $m=6$ and $m=4$ modes, respectively, as opposed to the preferred hydrodynamic onset mode of $m=7$. This saturation amplitude is almost independent of $\rm Pm$}, whilst dynamo onset is not, and so, whilst dynamo onset can occur almost arbitrarily close to ${\rm Ra}^{\prime}=1$ by increasing $\rm Pm$}, the saturated magnetic field strength should tend to zero in that limit. However, it has also been hypothesised that a subcritical dynamo, in which dynamo action occurs at Rayleigh numbers below hydrodynamic onset, could be possible since the onset of convection can be decreased by a sufficiently large magnetic field. Fearn Fearn (1979) finds that for asymptotically small $\rm Ek$},
$p_{10}\sim B_{0}\gtrsim\mathcal{O}\left({\rm Ek}^{1/6}\right)$ (23)
should be large enough to begin to decrease ${\rm Ra}_{c}$. Mason et al. Mason et al. (2022) find that ${\rm Ek}=10^{-4}$ is not sufficiently small to be asymptotic and that a larger field, $B_{0}=\mathcal{O}\left(1\right)$ is required to alter ${\rm Ra}_{c}$ (and also depends on $\rm Pm$}); these scalings are not amenable to the search for a subcritical dynamo.
@@
As the Rayleigh number increases (${\rm Ra}^{\prime}\gtrsim 5$), inertia grows and convective rolls become unstable, which characterises the onset of inertia-influenced solutions; these solutions can also sustain a dipolar magnetic field (for sufficiently large $\rm Pm$}), although the Lorentz force does not typically play a significant role, at least at low driving. At larger $\rm Ra$$^{\prime}$, a separate multipolar branch materialises due to the smaller inertial(/Coriolis) scales (compared with the (VAC) onset scales). However, a magnetic field can overcome resistance to large scales perpendicular to the rotation axis, generally increasing length scales on the strong-field (MAC) branch. The larger length scales on this branch reduce the energy lost to diffusion and hence increase the strength of the dipolar field; this feedback mechanism is another potential avenue for bistability.
The energy in zonal flow, which is on par with the other two components in hydrodynamic convection, is found to drop significantly in dynamo simulations, since it is resisted by the presence of a large scale poloidal field (Aubert, 2005). We investigate this aspect of the problem in §4, in our range of magnetoconvection simulations. In the absence of a significant $\Omega$-effect from a zonal flow, the dynamo mechanism on all three branches must be of an $\alpha^{2}$-type. A straightforward balance shows that the toroidal component of the field should therefore be of comparable strength to the poloidal field (Roberts and Soward, 1992). Teed & Dormy Teed and Dormy (2025b) verify this in the strong-field regime, but find that the toroidal field can be several times larger in the weak-field regime; a feature that can perhaps be explained by the disparity between the strengths of the axial and columnar flows in that case.
@@
Figure 4 shows the columnarity and symmetry in dynamo simulations for $\rm Ra$$^{\prime}$ between the onset of dynamo action and ${\rm Ra}^{\prime}=5$ for the two values of $\rm Pm$} that exhibit the clearest bistability between weak- and strong-field (dipolar) solutions. The sharp drops in equatorial symmetry at ${\rm Ra}^{\prime}=3.3$, for ${\rm Pm}=5$, and ${\rm Ra}^{\prime}=2.1$, for ${\rm Pm}=12$, indicate that an important element in the onset of strong-field convection is the breaking of equatorial symmetry which, as in hydrodynamic convection, seems to occur due to the destabilisation of antisymmetric modes of convection. Polar convective modes are significantly destabilised by the presence of an axial magnetic field since the wavelength of the onset mode perpendicular to the rotation axis is short, minimising the effect of rotation, which is felt more strongly inside the tangent cylinder; this also corresponds to the form required to minimise the bending of axial field lines (Sakuraba, 2002). The columnarity drops at the onset of the polar modes due to the growth of the magnetic field which destabilises convective rolls in the flow outside the tangent cylinder.
@@ 4.1 Magnetoconvection with ${\rm Pm}=1$
Figure 6: Time-averaged compensated-curl of forces (see Teed and Dormy, 2023) in simulations at ${\rm Pm}=1$, $p_{10}=0.1,1.0$, at three values of $\rm Ra$$^{\prime}$. Odd and even $l$-modes are presented separately in the first two cases due to the strong equatorial symmetry of the buoyancy force and the prominence of the $m=8$ convective mode and $m=0$ zonal mode leading to a dearth of odd-$l$ buoyancy modes.
@@ 4.2 Magnetoconvection with ${\rm Pm}=5$
Figure 10: Force spectra as in Figure 6, at ${\rm Pm}=5$, for ${\rm Ra}^{\prime}=2.5$ and $5.0$ for $p_{10}=0.1$ and ${\rm Ra}^{\prime}=1.6,1.8,2.5$, and $5.0$ for $p_{10}=1.0$.
@@
Figures 10(a),10(b), and10(d) show the force balances at ${\rm Ra}^{\prime}=1.6,\ 1.8,\ 2.5$, below, within, and above the vacillating regime, respectively. In each case the force balance could be described as MAC, as the Lorentz force is consistently part of the dominant balance across a range of scales, however it is clear that there are subtle distinctions between the three cases. At ${\rm Ra}^{\prime}=1.6$ and ${\rm Ra}^{\prime}=1.8$, the Lorentz force does not enter the dominant balance at the largest scales, entering sharply at $l=8$ for ${\rm Ra}^{\prime}=1.6$, when the columns are steady, and gradually at $l\approx 10$, for ${\rm Ra}^{\prime}=1.8$, when the columns are vacillating The remaining forces are also smoother at ${\rm Ra}^{\prime}=1.8$ than at ${\rm Ra}^{\prime}=1.6$, consistent with the wide range of scales that emerge due to the break-up of the columns. Above the vacillating regime the Lorentz force also enters the dominant balance at smallest $l$, and viscosity and inertia are increasing subdominant.
@@ 4.3 Magnetoconvection with ${\rm Pm}=12$
Figure 19(a) shows the zonally averaged poloidal streamlines for a weak-field solution at a bistable point in parameter space. The streamlines are equatorially symmetric and primarily reflect the axial circulation within convection columns, with neighbouring counter-rotating cells persisting after zonal averaging due to slight asymmetries between cyclonic and anti-cyclonic columns.
On the strong-field branch, the zonally averaged streamlines remain largely equatorially symmetric (Figure 19(d)), consistent with $s_{u}>0$ (Figure 5(f)), but the circulation cells extend much further into the equatorial region and inside the tangent cylinder. This reflects the increased magnetic field strength, which allows the Lorentz force to overcome the rotational constraint in regions where the boundaries are closest and (hence) the Coriolis force is most strongly felt.
@@ 5 Conclusions
In §3.2 we find that the weak-strong transition in dynamo simulations corresponds to a spontaneous breaking of the equatorial symmetry of the weak-field, highly-columnar, state as the Rayleigh number is increased, the abruptness of which indicates the presence of a secondary instability. There are three plausible candidates:
@@
    That a new convective instability occurs, possibly triggered by, but evolving independently of, the existing convection rolls.
Case (i) seems plausible since the inertia of convection rolls increases with the Rayleigh number. At sufficiently large $\rm Ra$} columnar flow can then become unstable to a buckling instability (like the Billant instability), which could lead to the emergence of modes, antisymmetric about the equator. However, this hypothesis is not supported by §3.1, and Figures 1(a),1(b), in particular, which show that the columnarity of the flow (in most of the shell) is reduced at lower Rayleigh numbers than the appearance of symmetry-breaking, which occurs at ${\rm Ra}^{\prime}\approx 10$. By this point, the columnarity of the flow has fallen to around $0.6$, which is significantly smaller than in some (weakly supercritical) strong-field states (Figures 4(a),4(b)). Nevertheless, this instability may be responsible for an alternative transition (at these large Ra) to inertially-driven/inertially-influenced states that may correspond/be relevant to multipolar solutions commonly found in dynamo simulations.
@@
Gilman Gilman (1975) (also Sreenivasan and Jones (2006)) has previously considered the consequences of the onset of convection inside the tangent cylinder, which can be well-approximated by the linear instability of a magnetic conduction state due to the relative quiescence of the tangent cylinder as a result the enhanced constraint of rapid rotation due to the closeness, in $z$, of the inner and outer boundaries. The Lorentz force, in rapidly-rotating magnetoconvection, can overcome the Taylor-Proudman constraint at significantly lower Rayleigh numbers than the viscous force alone (Fearn, 1979). Sreenivasan & Jones Sreenivasan and Jones (2006) find that an axial magnetic field reduces the critical Rayleigh number by around a factor of three, and the further addition of a strong toroidal magnetic field, which is also present in the tangent cylinder in models of weak-field convection (Figure 14), would only compound this effect. At ${\rm Pm}=5$, symmetry-breaking can occur at around ${\rm Ra}^{\prime}=3.3$, (three times less than in purely hydrodynamic simulations where symmetry-breaking must wait for inertial effects), whilst at ${\rm Pm}=12$, this occurs at ${\rm Ra}^{\prime}=2$.
Appendix B outlines a possible reasoning for polar convection to appear preferentially as an antisymmetric, rather than an symmetric mode, despite the disconnection of the regions within the tangent cylinder above and below the inner core. Hence, by measuring the departure of the flow from equatorial symmetry, we may capture the onset of polar convection.
Untangling the cause and effect in the correlation between the onset of polar convection and the onset of the strong-field regime is challenging and could be an interesting subject for future work. Skene et al. Skene et al. (2024) investigate the case of bistability between a hydrodynamic and a weak-field state using an adjoint method, and hence are able to determine an unstable fixed point that influences the evolution of an initial state towards either a successful or failed dynamo. If similar methods can be employed between the weak- and strong-field branches it may help to understand the exact role that the onset of polar convection has in the development of the strong-field regime.
@@ Appendix A Magnetic boundary conditions
We enforce that the radial component of the magnetic field, $\boldsymbol{\hat{r}}\cdot\boldsymbol{B}$, is fixed at the outer boundary, whilst the inner boundary corresponds to electrically-insulating magnetic boundaries. These are aphysical conditions, but chosen to promote desirable attributes of simulations. We detail these conditions, alongside other possible magnetoconvection conditions below, for comparison.
Electrically-insulating boundaries
To a reasonable approximation, the Earth’s mantle is electrically insulating. In an insulating medium, the current vanishes,
@@
Since the magnetic field and the current (its curl) are both solenoidal fields, they can be decomposed into poloidal and toroidal components as
@@
$\boldsymbol{j}$ $=\boldsymbol{\nabla}\times\boldsymbol{\nabla}\times\left(j_{p}\boldsymbol{r}\right)+\boldsymbol{\nabla}\times\left(j_{t}\boldsymbol{r}\right),$ (26b)
respectively. By definition, $\boldsymbol{B}$ and $\boldsymbol{j}$ are related by
@@
and hence, the magnetic field in an insulating medium must satisfy
@@
$B_{p}=\sum_{l}\sum_{m}P^{m}_{l}\left(r\right)Y^{m}_{l}\left(\theta,\phi\right),\quad B_{t}=\sum_{l}\sum_{m}T^{m}_{l}\left(r\right)Y^{m}_{l}\left(\theta,\phi\right),$ (29)
@@
If there are no external sources then the exterior magnetic field is produced entirely by fluid in the spherical shell and $P^{m}_{l}=Ar^{-l-1}$, for some constant $A$; the boundary condition on the internal field, matching with this solution is
$\left(r\frac{{\rm d}}{{\rm d}r}+l\left(l+1\right)\right)P^{m}_{l}\biggr|_{r=r_{\rm o}}=0,\quad\forall\penalty 10000\ l,m.$ (31)
At the inner boundary the role of internal and external sources is reversed by the geometry in relation to increasing $r$, i.e., sources correspond to $\alpha=l$, therefore, the no-source condition at the inner boundary is
@@
External sources
Say an external source at the outer boundary induces a uniform axial magnetic field of the form $\boldsymbol{B}=B_{0}\boldsymbol{\hat{z}}$ where $\boldsymbol{\hat{z}}=\cos{\theta}\boldsymbol{\hat{r}}-\sin{\theta}\boldsymbol{\hat{\theta}}$. Then, given $B_{r}=r^{-1}L^{2}B_{p}$ (where $L^{2}=l(l+1)$, is the angular component of the spherical Laplacian), the axial field is
$B_{p}(r,\theta,\phi)=P^{0}_{1}(r)\cos{\theta},\quad P^{0}_{1}(r)=\frac{B_{0}}{2}r.$ (33)
@@
$P^{0}_{1}=\frac{B_{0}}{2}r+\frac{A^{0}_{1}}{r^{2}},\qquad P^{m}_{l}=A^{m}_{l}r^{-l-1},\penalty 10000\ \left(l,m\right)\neq\left(1,0\right),$ (34)
@@
$\left(\frac{{\rm d}}{{\rm d}r}+\frac{2}{r}\right)P_{1}^{0}(r)\biggr|_{r=r_{\rm o}}=\frac{3B_{0}}{2},$ (35)
@@
This form of magnetoconvection is adopted by Mason et al. (2022), for example.
Fixed radial magnetic field
A significant difference between weak- and strong-field dynamos is the strength of the dipolar component of the magnetic field. To better understand the direct effect this has on the flow, we impose a fixed morphology and strength at the outer boundary. To match with the electrically-insulating conditions above, we maintain the zero condition on the toroidal scalar of the magnetic field as zero but fix the poloidal scalar, i.e., the radial component of the field.
Say that we set the dipolar component of the poloidal function, $P_{1}^{0}(r_{\rm o})=C$, then the radial component of the magnetic field is given by
@@
and hence, with all other $P^{m}_{l}=0$, $\left(l,m\right)\neq\left(1,0\right)$,
$B_{r}\left(r_{\rm o},\theta,\phi\right)=\frac{2C}{r_{\rm o}}\cos{\theta}=\frac{13C}{10}\cos{\theta},$ (37)
where $r_{\rm o}=20/13$, when using radius ratio, $\chi=0.35$, and normalised shell thickness. Fixing the poloidal scalar at the boundary is equivalent to perfectly electrically-conducting conditions on the poloidal component of the magnetic field, but not on the toroidal component, where $T_{l}^{m}=0$ from the insulating conditions is retained.